# Stronger Than You Think: Benchmarking Weak Supervision on Realistic Tasks

### v1

@@ Abstract
Weak supervision (WS) is a popular approach for label-efficient learning, leveraging diverse sources of noisy but inexpensive weak labels to automatically annotate training data. Despite its wide usage, WS and its practical value are challenging to benchmark due to the many knobs in its setup, including: data sources, labeling functions (LFs), aggregation techniques (called label models), and end model pipelines. Existing evaluation suites tend to be limited, focusing on particular components or specialized use cases. Moreover, they often involve simplistic benchmark tasks or de-facto LF sets that are suboptimally written, producing insights that may not generalize to real-world settings. We address these limitations by introducing a new benchmark, BoxWRENCH, [^1] The box wrench is the most ubiquitous and practical wrench. designed to more accurately reflect real-world usages of WS. This benchmark features tasks with (1) higher class cardinality and imbalance, (2) notable domain expertise requirements, and (3) multilingual variations across parallel corpora. For all tasks, LFs are written using a careful procedure aimed at mimicking real-world settings. In contrast to existing WS benchmarks, we show that supervised learning requires substantial amounts (1000+) of labeled examples to match WS in many settings.
@@ 1 Introduction
WS frameworks typically have a simple three-stage approach. First, they formalize weak sources into labeling functions (LFs). In contrast to manual labeling, these can be automatically applied to an entire unlabeled dataset. Next, since LFs are inherently noisy and may conflict with one another, a label model (LM) is used to estimate the quality of each source (typically without access ground truth labels) and then to aggregate their outputs into high-quality pseudolabels. Finally, these pseudolabels can be used to train a downstream model. A vast literature studies variations on this basic recipe, with diverse approaches to crafting LFs, creating LMs, and noise-aware training of end-models [45].
For practitioners, a key question is when is WS useful? While it is natural to produce benchmarks that answer this, surprisingly, there has been relatively little work doing so. One reason for this may be the overall complexity of WS pipelines. The performance of a WS system varies with (1) the underlying task and data, (2) the LFs, (3) the choice of LM, and (4) the choice of end model and training procedure. Several benchmarks predominantly focus on only one of these. For example, WRENCH [44] focuses on evaluating (3), the LM, while AutoWS-Bench-101 [35] focuses primarily on (2), the LFs, and specifically, techniques for automatically generating model-based LFs.
Recently, Zhu et al. [47] tackles the goal of quantifying the value of WS. They argue that the benefits of WS are often overestimated by showing that fine-tuning on only 50 ground-truth labels can achieve comparable—or better—results than certain WS approaches for many benchmark datasets. They suggest that WS may not be broadly useful, as obtaining 50 “clean” labels is rarely prohibitive, and data at this scale (or larger) may still be needed for tuning or evaluation even when using WS.
In this work, we show that these findings result from the simplicity of existing datasets rather than the inherent weakness of WS. In particular, we identify two main issues with the current WS benchmarks that led to this result, and show that WS may be stronger than is thought in more realistic settings:
@@
We introduce a new benchmark, BoxWRENCH, that addresses these two challenges. It enables us to quantify the practical advantages of WS in a wide range of settings. Our findings indicate that even simple WS approaches often provide substantial value. We address the issues we identified by
@@
  • Showing that by adhering to careful LF design practices, we can write effective LFs for these tasks that can even improve upon existing benchmark LFs.
Additionally, we experiment with the re-usability of LFs across related task specifications, another key practical advantage of WS deployments that does not show up in standard benchmarks. Using the MASSIVE dataset [10], we introduce a new method to reuse existing LFs on multi-language datasets, providing an out-of-the-box performance boost.
Our benchmark consists of five text-classification WS tasks that showcase the power of WS in a variety of challenging real-world scenarios. For two of our tasks, we produce new LFs, while for one, we improve the existing LFs from WRENCH [44]. The design of these LFs follows a rigorous procedure that we release as part of our benchmark, acting as guidance for LF design and for WS benchmarking overall. We publicly release all experiment code and datasets. [^2] https://github.com/jeffreywpli/stronger-than-you-think
@@ 2 Background and Related Work
We provide a brief background on WS techniques and benchmarking efforts. We note in passing that the term WS can be overloaded and is also applied to other families of techniques, particularly in computer vision [27, 20].
Here, we use WS to refer to approaches that fall under programmatic WS [32, 45].
Weak Supervision. WS uses imperfect label sources, often small programs, as LFs to generate labels for unlabeled data. Perhaps the most popular types of LFs are heuristics or rules obtained from subject matter experts [31] encoded into programs. Other potential sources include knowledge base queries, pretrained models, and more [16, 21, 8, 28, 17]. Many works focus on either improving the aggregation technique (the LM) or obtaining improved approaches to crafting LFs. Other methods aim to improve the end model learning algorithms when training on WS pseudo-labels (e.g. COSINE [43]). The simplest aggregation technique is directly using majority vote. Others seek to learn (without ground truth labels) the accuracy of the LFs and thus perform a higher-quality aggregation [31, 13, 38]. For LF construction, variations such as learning small models on tiny amounts of data [41], or using code-generating large language models to craft LFs have been proposed [19, 18, 15].
@@
Existing WS benchmarks typically focus on a particular component of a WS system or a particular use case. WRENCH [44] benchmarks LMs and is therefore aimed at aggregation. AutoWS-Bench-101 [35], in contrast, studies the effectiveness of automated LF construction techniques. Finally, WALNUT [46] studies WS techniques in the context of natural language understanding. All of these benchmarks are highly useful, but do not attempt to measure the value of WS techniques more broadly. A recent effort by Zhu et al. [47] tackles this question and finds that in particular settings, WS may not be of great value. Specifically, it suggests that only a small amount of labeled data is sufficient to train a supervised model to a level of quality equivalent to that provided by WS. We are inspired by this work, studying whether we can obtain similar findings across a broader range of realistic scenarios.
@@ 3.1 Goals
The ultimate goal of BoxWRENCH is to bridge WS research and practice by introducing more realistic benchmarks for WS. The first step towards such a goal is to gain a better understanding of the question: when is WS useful? To do so, we first gather a suite of datasets which addresses two key areas in which current WS benchmarks fall short.
@@
    Benchmark datasets tend be simplistic, exhibiting properties that are not representative of many real-world problems. This includes having a small label space (often binary), balanced label distributions, and relying on general rather than domain-specific knowledge.
@@
    Current WS benchmarks are used with de-facto LF sets that vary in quality. A poorly written LF set may also result in a less realistic benchmark (e.g., if a task involves domain expertise but experts were not involved in writing the LFs).
To address the first issue, we introduce WS tasks that directly target the aforementioned gaps: focusing on those with greater class counts, class imbalance, and domain-specificity. To address the second issue, we place care into writing higher-quality LFs for all datasets, including improving existing LFs. Using these datasets, we aim to measure how many labeled examples are needed before supervised learning catches up to WS techniques. We formalize this notion by plotting the performance curves of WS techniques and fully supervised learning (as functions of the number of labels) and analyzing the corssover points where these curves intersect. To show that WS is effective, the crossover point in which supervised learning surpasses WS should be sufficiently high, i.e., WS cannot be matched by simply labeling a trivial amount of examples. Using crossover points to measure the effectiveness of WS, we aim to establish a regime in which WS is practically useful on our suite of more challenging and realistic datasets.
@@ 3.2 Problem Formulation
Let $\mathcal{D}\subseteq\mathcal{X}\times\mathcal{Y}$ be our data distribution. We first sample an unlabeled training set with $n$ examples $X_{\text{train}}=[x_{i}]_{i=1}^{n}$, and a small labeled validation set of $m$ examples: $D_{\text{val}}=[X_{\text{val}},Y_{\text{val}}]$ with $X_{\text{val}}=[x_{i}]_{i=n+1}^{n+m}$ and $Y_{\text{val}}=[y_{i}]_{i=n+1}^{n+m}$ with $x_{i}\in\mathcal{X}$ and $y_{i}\in\mathcal{Y}$. We are interested in learning a function $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$ that minimizes the expected risk $R(f_{\theta})=\mathbb{E}_{(x,y)\sim\mathcal{D}}[L(f_{\theta}(x),y)]$ where $L$ is a loss function and $\theta$ are the model parameters. We then evaluate $f_{\theta}$ on an i.i.d. drawn test set, $D_{\text{test}}$. In the case of WS, let $\mathcal{H}$ be the hypothesis class of the LFs. For every $h\in\mathcal{H}$, we have $h:\mathcal{X}\rightarrow\mathcal{Y}\cup\{-1\}$, where $-1$ denotes an abstention. Let $\text{LF}_{\text{ws}}\subset\mathcal{H}$ be the set of LFs that we use to generate LF votes. For every $x_{i}\in X_{\text{train}}$, $\lambda_{i,j}=h_{j}(x_{i})$, where $h_{j}\in\text{LF}_{\text{ws}}$, is $j$-th LF votes for $x_{i}$. Using an LM that aggregates the LF votes, we obtain $\widehat{y}_{i}=\text{LM}\left([\lambda_{i,j}]_{j\in[|\text{LF}_{\text{WS}}|]}\right)$, namely, the weak label for $x_{i}$. We construct the weak labels for the training set as $\widehat{Y}_{\text{train}}=[\widehat{y}_{i}]_{i=1}^{n}$. We then use $D_{\text{WS}}=[X_{\text{train}},\widehat{Y}_{\text{train}}]$ as the weakly labeled training set, to train a model $f_{\theta_{\text{WS}}}$.
Following the continuous fine-tuning (CFT) setting of Zhu et al. [47], we start with a pre-trained model $f_{\theta_{\text{WS}}}$ and fine-tune it on the clean validation set $D_{\text{val}}$ to obtain $f_{\theta_{\text{CFT}}}$. An important baseline is a supervised-only approach, where we train the model directly on labels from the clean validation set $D_{\text{val}}$ to obtain $f_{\theta_{\text{SUP}}}$.
@@ 3.3 Datasets
Existing WS benchmarks often exhibit low class cardinality (under 5) or use label distributions that are highly balanced (i.e. near uniform). However, real-world WS tasks often do not conform to these assumptions, as they often involve imbalanced label distributions and high class cardinality. To evaluate WS in realistic regimes, we propose a collection of datasets [^3] With the current exception of Amazon31, which was recently retracted, all of the datasets we use are publicly available and with proper licenses (see Appendix B) with varying levels of difficulty and requirements for domain expertise, having a range of class distributions and cardinalities. We first describe each of the datasets used in BoxWRENCH, with their metadata shown in Table 1, and describe our LF design procedure.
@@
  • ChemProt [22] is a chemical relation classification dataset comprising 1,820 PubMed abstracts with chemical-protein interactions annotated by domain experts. The dataset was studied in [2, 47]. Previous works on WS created LFs for the dataset and showed the efficacy of WS in the dataset [43]. We push the boundaries of WS in the dataset by modifying the LFs to incorporate the distance between the chemical and protein entities in the text, among other minor modifications (see Fig. 13).
  • Claude9444Claude9 is imbalanced, with 90%+ of data belonging to one class, so we evaluate using macro-averaged F1. is based on UNFAIR-ToS [7, 25], which includes 50 Terms of Service (ToS) from online platforms and sentence-level annotations with 8 types of unfair contractual terms (potentially violating user rights according to EU consumer law). [^5] Art.3 of Direct. 93/13, Unfair Terms in Consumer Contracts (http://data.europa.eu/eli/dir/1993/13/oj). LFs for this dataset were created by a law graduate student.
  • MASSIVE18 [10, 23] is a parallel corpus of 52 languages with annotations for Natural Language Understanding tasks of intent prediction and slot annotation. The dataset is parallel, because given any two languages subsets, there exist a bijection between the examples in the two subsets. We construct several WS datasets based on the original MASSIVE, we also provide pipeline that can generate similar datasets in different languages easily.
    • MASSIVE60 [10, 23] we created a high cardinality version of MASSIVE18 that has weak labels by splitting existing classes of MASSIVE18 into detailed sub classes. The original LFs have been generalized, with a mechanism that selects a random subset from the original results for each instance (see Fig. 9).
  • Amazon31 is built from the Amazon product reviews [1] dataset consisting of reviews and their categories. [^6] This dataset has been taken down. We include it in experiments but we will not release it in BoxWRENCH. Due to each class’s high overlap and conflict rate, we merged several labels and reduced the class cardinality to 31.
@@
We randomly selected samples with clean labels from the training set to create a development set for LFs. We use 250 examples as a development set for Amazon31, which is consistent with the development of LFs for Banking77 and MASSIVE18 in [23]. For Claude9, the development set had 24 examples. We manually inspected labeled examples in the development set, and identified patterns for each class. Then we created multiple keyword, dictionary-based, and regular expression-based LFs for each class. We calculated LF statistics (which do not require label information) on the original training and validation sets, including coverage and LF conflict ratios. To to evaluate the final LFs, we calculate their accuracy scores on the original validation set. [^7] We created the LFs ourselves and did not hire annotators.
Figure 1: We compare end models trained with three different pipelines: (1) Supervised, which fine-tunesuses clean labels from the validation set for fine-tuning; (2) Weakly Supervised, which applies all LFs and aggregates them with the LM then fine-tuning an EM using the training data with the aggregated labels; using data with clean labels from the validation set for hyperparameter searching or early stopping. Continuous-Fine-Tuning: using parameters of the EMs after WS experiments, and then continuously fine-tuning them on the same data with clean validation labels.

### v2

@@ Abstract
Weak supervision (WS) is a popular approach for label-efficient learning, leveraging diverse sources of noisy but inexpensive weak labels to automatically annotate training data. Despite its wide usage, WS and its practical value are challenging to benchmark due to the many knobs in its setup, including: data sources, labeling functions (LFs), aggregation techniques (called label models), and end model pipelines. Existing evaluation suites tend to be limited, focusing on particular components or specialized use cases. Moreover, they often involve simplistic benchmark tasks or de-facto LF sets that are suboptimally written, producing insights that may not generalize to real-world settings. We address these limitations by introducing a new benchmark, BoxWRENCH, [^1] The box wrench is the most ubiquitous and practical wrench. designed to more accurately reflect real-world usages of WS. This benchmark features tasks with (1) higher class cardinality and imbalance, (2) notable domain expertise requirements, and (3) opportunities to re-use LFs across parallel multilingual corpora. For all tasks, LFs are written using a careful procedure aimed at mimicking real-world settings. In contrast to existing WS benchmarks, we show that supervised learning requires substantial amounts (1000+) of labeled examples to match WS in many settings.
@@ 1 Introduction
WS frameworks typically have a simple three-stage approach. First, they formalize weak sources into labeling functions (LFs). In contrast to manual labeling, these can be automatically applied to an entire unlabeled dataset. Next, since LFs are inherently noisy and may conflict with one another, a label model (LM) is used to estimate the quality of each source (typically without accessing ground truth labels) and then to aggregate their outputs into high-quality pseudolabels. Finally, these pseudolabels can be used to train a downstream model. A vast literature studies variations on this basic recipe, with diverse approaches to crafting LFs, creating LMs, and noise-aware training of end-models [47].
For practitioners, a key question is when is WS useful? While it is natural to produce benchmarks that answer this, surprisingly, there has been relatively little work doing so. One reason for this may be the overall complexity of WS pipelines. The performance of a WS system varies with (1) the underlying task and data, (2) the LFs, (3) the choice of LM, and (4) the choice of end model and training procedure. Several benchmarks predominantly focus on only one of these. For example, WRENCH [46] focuses primarily on evaluating (3), the LM, while AutoWS-Bench-101 [37] focuses on (2), the LFs, and specifically, techniques for automatically generating model-based LFs.
Recently, Zhu et al. [49] tackle the goal of quantifying the value of WS. They argue that the benefits of WS are often overestimated by showing that fine-tuning on only 50 ground-truth labels can achieve comparable—or better—results than certain WS approaches for many benchmark datasets. They suggest that WS may not be broadly useful, as obtaining 50 “clean” labels is rarely prohibitive, and data at this scale (or larger) may still be needed for tuning or evaluation even when using WS.
In this work, we show that these findings result from the simplicity of existing datasets rather than the inherent weakness of WS. In particular, we identify key issues with the current WS benchmarks that led to this result and show that WS may be stronger than is thought in more realistic settings:
@@
  3. 3.
    Previous benchmarks do not capture the adaptability of LFs across task specification, a key practical advantage of WS deployments compared to manual labeling.
We introduce a new benchmark, BoxWRENCH, that addresses these three challenges. It enables us to quantify the practical advantages of WS in a wide range of settings. Our findings indicate that even simple WS approaches often provide substantial value. We address the issues we identified by:
@@
  • Showing that by adhering to careful LFs design practices, we can write effective LFs for these tasks that can even improve upon existing benchmark LFs.
  • Using the MASSIVE dataset [10], we study simple but effective strategies for adapting existing LFs written for English data to parallel versions of the task in other languages.
Our benchmark consists of five text-classification WS tasks that showcase the power of WS in a variety of challenging real-world scenarios. For two of our tasks, we produce new LFs, while for one, we improve the existing LFs from WRENCH [46]. The design of these LFs follows a rigorous procedure that we release as part of our benchmark, acting as guidance for LF design and for WS benchmarking overall. We publicly release the code and other assets for our study at https://github.com/jeffreywpli/stronger-than-you-think.
@@ 2 Background and Related Work
We provide a brief background on WS techniques and benchmarking efforts. We note that the term WS can be overloaded, as it is also applied to other families of techniques that generally aim to learn from indirect or noisy forms of supervision [32], particularly in computer vision [28, 20]. Here, we use WS to refer to approaches that fall under programmatic WS [34, 47].
Weak Supervision. WS aggregates multiple imperfect label sources, each formalized as a labeling function (LF), to synthesize labels for unlabeled data. Perhaps the most popular types of LFs are heuristics or rules obtained from subject matter experts [33] encoded into programs. Other potential sources include knowledge-base queries, pretrained models, and more [16, 21, 8, 29, 17]. Many works in WS focus on improving either how LFs are crafted or how they are aggregated. For LF construction, variations such as learning small models on tiny amounts of data [43], or using code-generating large language models to craft LFs have been proposed [19, 18, 15]. For aggregation, the simplest technique is to perform a majority vote while other methods seek to infer (without using ground truth labels) the accuracy of the LFs with a label model (LM) and thus perform a higher-quality aggregation [33, 13, 40]. Orthogonal to both directions, other works study more effective strategies for training end models on WS-generated pseudo-labels [45, 24].
@@
Existing WS benchmarks typically focus on a particular component of a WS system or a particular use case. WRENCH [46] primarily benchmarks different label models and is therefore aimed at aggregation. AutoWS-Bench-101 [37], in contrast, studies the effectiveness of automated LF construction techniques. Finally, WALNUT [48] studies WS techniques in the context of natural language understanding. All of these benchmarks are highly useful, but do not attempt to measure the value of WS techniques more broadly. A recent effort by Zhu et al. [49] tackles this question and finds that in particular settings, WS may not be of great value. Specifically, it suggests that only a small amount of labeled data is sufficient to train a supervised model to a level of quality equivalent to that provided by WS. We are inspired by this work, studying whether we can obtain similar findings across a broader range of realistic scenarios.
@@ 3.1 Goals
The ultimate goal of BoxWRENCH is to bridge WS research and practice by introducing more realistic benchmarks for WS. The first step towards such a goal is to gain a better understanding of the question: when is WS useful? To do so, we first gather a suite of datasets that addresses two key areas in which current WS benchmarks fall short.
@@
    Benchmark datasets tend to be simplistic, exhibiting properties not representative of many real-world problems. This includes having a small label space (often binary), balanced label distributions, and relying on general rather than domain-specific knowledge.
@@
    Current WS benchmarks are used with de-facto LF sets that vary in quality. A poorly written LF set may also result in a less realistic benchmark (e.g. if a task involves domain expertise but experts were not involved in writing the LFs).
To address the first issue, we introduce WS tasks that directly target the aforementioned gaps: focusing on those with greater class counts, class imbalance, and domain-specificity. To address the second issue, we place care into writing higher-quality LFs for all datasets, including improving existing LFs. Using these datasets, we aim to measure how many labeled examples are needed before supervised learning catches up to WS techniques. We formalize this notion by plotting the performance curves of WS techniques and fully supervised learning (as functions of the number of labels) and analyzing the crossover points where these curves intersect. To show that WS is effective, the crossover point in which supervised learning surpasses WS should be sufficiently high, i.e., WS cannot be matched by simply labeling a trivial amount of examples. Using crossover points to measure the effectiveness of WS, we aim to establish a regime in which WS is practically useful on our suite of more challenging and realistic datasets.
@@ 3.2 Problem Formulation
Let $\mathcal{D}\subseteq\mathcal{X}\times\mathcal{Y}$ be our data distribution. We first sample an unlabeled training set with $n_{\text{train}}$ examples $X_{\text{train}}=[x_{i}]_{i=1}^{n_{\text{train}}}$, and a small labeled validation set of $n_{\text{val}}$ examples: $D_{\text{val}}=[X_{\text{val}},Y_{\text{val}}]$ with $X_{\text{val}}=[x_{i}]_{i=n_{\text{train}}+1}^{n_{\text{train}}+n_{\text{val}}}$ and $Y_{\text{val}}=[y_{i}]_{i=n_{\text{train}}+1}^{n_{\text{train}}+n_{\text{val}}}$ with $x_{i}\in\mathcal{X}$ and $y_{i}\in\mathcal{Y}$. We are interested in learning a function $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$ that minimizes the expected risk $R(f_{\theta})=\mathbb{E}_{(x,y)\sim\mathcal{D}}[L(f_{\theta}(x),y)]$ where $L$ is a loss function and $\theta$ are the model parameters. To estimate $R(f_{\theta})$, we assume the existence of an i.i.d. test set $D_{\text{test}}$, for which labels are also available. In the case of WS, we assume that labels for $X_{\text{train}}$ are provided by a set of labeling functions $\text{LF}_{\text{ws}}=\{\lambda_{j}\}_{j=1}^{m}$ where each LF, $\lambda_{j}:X\rightarrow Y\cup\{-1\}$ encodes some heuristic that labels or abstains (denoted by -1) on each input $x_{i}$. Using a label model LM that aggregates the LF votes, $\lambda_{i,j}=\lambda_{j}(x_{i})$, we obtain $\widehat{y}_{i}=\text{LM}\left([\lambda_{i,j}]_{j=1}^{m}\right)$, namely, the weak label for $x_{i}$. We construct the weak labels for the training set as $\widehat{Y}_{\text{train}}=[\widehat{y}_{i}]_{i=1}^{n_{\text{train}}}$. We then use $D_{\text{WS}}=[X_{\text{train}},\widehat{Y}_{\text{train}}]$ as the weakly labeled training set, to train a model $f_{\theta_{\text{WS}}}$.
Following Zhu et al. [49], we compare the performance $f_{\theta_{\text{WS}}}$ with two other models that train directly on the validation labels: $f_{\theta_{\text{SUP}}}$ is the model trained only on $D_{\text{val}}$, which serves as a critical baseline for assessing the usefulness of WS. $f_{\theta_{\text{CFT}}}$ is a model that results from using $f_{\theta_{\text{WS}}}$ as an initialization and further performs continuous fine-tuning on $D_{\text{val}}$. Weak supervision is considered useful given $D_{\text{val}}$ if the performance of either $f_{\theta_{\text{WS}}}$ or $f_{\theta_{\text{CFT}}}$ remains significantly higher than that of $f_{\theta_{\text{SUP}}}$.
@@ 3.3 Datasets
Existing WS benchmarks often exhibit low class cardinality (under 5), highly balanced label distributions (i.e. near uniform), and simplistic underlying tasks that do not require domain-specific knowledge. However, real-world tasks often do not conform to these assumptions. Thus, to evaluate WS in realistic regimes, we propose a collection of datasets [^2] With the current exception of Amazon31, which was recently retracted, all of the datasets we use are publicly available and with proper licenses (see Appendix B). with varying levels of class cardinality and imbalance, as well as requirements for domain expertise. We describe each of the datasets used in BoxWRENCH, with their metadata shown in Table 1, and describe their LFs.
@@
  • ChemProt [22] is a chemical relation classification dataset comprising 1,820 PubMed abstracts with chemical-protein interactions annotated by domain experts. The dataset was studied in [2, 49]. Previous works on WS created LFs for the dataset and showed the efficacy of WS in the dataset [45]. We push the boundaries of WS in the dataset by modifying the LFs to incorporate the distance between the chemical and protein entities in the text, among other minor modifications (see Appendix C and Fig. 13).
  • Claude9333Claude9 is imbalanced, with 90%+ of data belonging to one class, so we evaluate using macro-averaged F1. is based on UNFAIR-ToS [7, 26], which includes 50 Terms of Service (ToS) from online platforms and sentence-level annotations with 8 types of unfair contractual terms (potentially violating user rights according to EU consumer law). [^4] Art.3 of Direct. 93/13, Unfair Terms in Consumer Contracts (http://data.europa.eu/eli/dir/1993/13/oj). LFs for this dataset were created by one of the authors of this work who is a law graduate student.
  • MASSIVE{18,60} [10, 23] is a corpus of human-to-voice assistant interactions, where the task is to predict one of 60 fine-grained or 18 coarse-grained annotations of user intents. The corpus includes parallel versions across 52 different languages, where each example is present in each language. In our work, we explore reusing the existing LFs written for the English variant of the 18-class prediction task to generalize to additional languages and more fine-grained classes. For non-English variants of MASSIVE18, we apply English LFs by first translating each example back to English using DeepL (see Fig. 2). For MASSIVE60, we leverage the original LFs from MASSIVE18 to predict the superclass for each instance. To generalize from coarse to fine-grained predictions in MASSIVE60, we randomly select a subclass corresponding to the predicted superclass from the original LFs.
  • Amazon31 is built from the Amazon product reviews [1] dataset consisting of reviews and their categories. [^5] This dataset has been taken down. We include it in experiments, but we will not release it in BoxWRENCH. Due to each class’s high overlap and conflict rate, we merged several labels and reduced the class cardinality to 31.
@@
We randomly selected samples with clean labels from the training set to create a development set for LFs. We use 250 examples as a development set for Amazon31, which is consistent with the development of LFs for Banking77 and MASSIVE18 in [23]. For Claude9, the development set had 24 examples. We manually inspected labeled examples in the development set, and identified patterns for each class. Then we create multiple keyword-based, dictionary-based, and regular expression-based LFs for each class [^6] We created the LFs ourselves and did not hire annotators.. We calculated LF statistics (which do not require label information) on the original training and validation sets, including coverage and LF conflict ratios. To evaluate the final LFs, we calculate their accuracy scores on the original validation set.
Figure 1: We compare end models trained with three different pipelines: (1) Supervised, which uses clean labels from the validation set for fine-tuning; (2) Weakly Supervised, which trains on labels obtained by applying all LFs and aggregating them with the label model, using the validation set for hyperparameter search and early stopping. Continuous-Fine-Tuning: which takes a weakly supervised model and then continuously fine-tuning it on the same clean validation labels.